OpenStudy (anonymous):

Compute ∫1 -1 (i don't know how to write this properly in open study) 2x cos(x2)dx

OpenStudy (badhi):

\[\int 1-2x\cos(x^2)\,dx\] is this your question?

OpenStudy (badhi):

if so, use the substitution\[u=x^2\] then, \[du=2xdx\] now back to the integral \[\int 1-2x\cos(x^2)\,dx=\int\,dx-\int \cos(u)\,du=x-\sin(u)+c=x-\sin(x^2)+c \]